# Ruishu Liu, Yulin Cao,  Xiaojie Wang

### v1

@@ 1. Case 1:
    $\kappa+1>2\rho$.
    Due to Lemma 2.2 and Lemma 3.6, one can infer that
@@
    Equipped with (4.11) as well as Lemmas 2.4, 2.6 one can directly obtain
@@ 2. Case 2:
    The difference for the critical case comes from the extra restriction that $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$, as required by Assumption 3.3, which together with Lemmas 2.3, 3.6 ensures the moment boundedness of $\mathbb{E}\big{[}|X_{t}|^{p}\big{]}$ and $\mathbb{E}\big{[}|Y_{n}|^{p}\big{]}$ for $p\leq 2\gamma+1\leq\tfrac{2c_{2}}{c_{3}^{2}}+1$. Therefore, on the condition $\tfrac{c_{2}}{c_{3}^{2}}>\kappa-\tfrac{1}{2}$, one can derive $2\kappa$-th moment bounds for both analytic and numerical solutions
@@
    Recalling (3.8) one knows that
@@
    by Corollary 3.7. This together with Lemma 2.5 and 2.6 helps us to deduce that, on the more strict assumption $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$,
@@
Let $\{X_{t}\}_{0\leq t\leq T}$ and $\{Y_{n}\}_{0\leq n\leq N}$ be defined as (2.2) and (3.1), respectively. Let $f_{h}$ and $g_{h}$ be defined as in Example 3.5. Suppose that Assumptions 3.2, 3.3 and 3.4 hold, with additionally one of the following conditions standing:
@@
Then there exists a constant $C>0$ independent of the step-size $h$ such that
@@
We would like to mention that, for the general critical case $\kappa+1=2\rho$ one can observe a slightly more strict restriction on the model parameters than that for the implicit methods (STMs) in [20], where $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$ is required. This is caused by the need of $m_{1},m_{2}\leq 2\gamma+1$ in order to confirm Assumption 3.4. More precisely, as indicated by Example 3.5, for the general critical case, $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ as well as $12(2\alpha+1)\rho=m_{2}<m_{1}=2(2\alpha+1)\kappa\leq 2\gamma+$ 1 are required, leading to $(2\alpha+1)\kappa\leq\tfrac{c_{2}}{c_{3}^{2}}+\tfrac{1}{2}$ as mentioned.
@@ 5 Numerical experiments
The semi-implicit tamed Euler-type method (TEM) in Example 3.5 with $\alpha=\tfrac{1}{2}$ is applied to approximate the the model:
@@
The following two sets of parameters are taken as numerical tests.
@@
$\kappa=2,\rho=1.5,c_{-1}=1.5,c_{0}=2,c_{1}=1,c_{2}=4,c_{3}=1$.
Note that Example 5.1 satisfies the non-critical condition $\kappa+1>2\rho$, while Example 5.2 satisfies the general critical condition $\kappa+1=2\rho$. The strong convergence rates are tested with fixed total time $T=1$. The numerical approximations with the step-size $h_{exact}=2^{-14}$ is used to simulate the ”exact” solutions, while $h=2^{-i},i=4,5,...,9$ are used for numerical solutions over $10^{4}$ Brownian paths. The estimation error is calculated in terms of $e_{h}:=\bigg{(}\sup\limits_{0\leq n\leq N}\mathbb{E}\Big{[}|X_{t_{n}}-Y_{n}|^{2}\Big{]}\bigg{)}^{\tfrac{1}{2}}$. The stochastic theta method (STM) [20] with $\theta=1$, which coincides with the backward Euler-Maruyama method (BEM) [18], is also tested in comparison with TEM.
Fig. 1 plots the average sample error against various step-sizes on a log-log scale, with BEM in a black solid line and TEM in a red solid line. It can be easily observed the order $0.5$ strong convergence for both two schemes. By a least-squares fitting, the resulting convergence rate $q$ and the least squares residual are shown in Table 1, where one can also confirm the predicted convergence rate.
Moreover, the computational costs of BEM and the proposed TEM scheme over $10^{4}$ Brownian paths are shown in Table 2 and 50 out of the $10^{4}$ paths are shown in Fig. 2. It is easy to see that the TEM can preserve positivity and reduce computational costs sigficantly, compared with BEM, especially as the parameter $\kappa$ increases.
Figure 1: Strong convergence rate of the semi-implicit taming Euler method under Example 5.1 (left) and Example 5.2 (right).
Figure 2: 50 paths of numerical solutions for BEM and TEM
Table 1: A least squares fit for the convergence rate $q$
@@
Case 1 $q=0.5409$, $resid=0.0269$ $q=0.5173$, $resid=0.0128$
Case 2 $q=0.5558$, $resid=0.0698$ $q=0.5288$, $resid=0.0326$
@@ 6 Conclusion
In this paper we propose and analyze an explicit time-stepping scheme for the strong approximations of the generalized Aït-Sahalia model. Based on a semi-implicit strategy and proper modifications, the numerical scheme is designed to be positivity preserving and easily implementable with cheap computational costs compared with existing positivity preserving schemes. The mean-square strong convergence rate of the scheme is carefully analyzed and proved to be order $0.5$. Numerical experiments are also provided to support the theoretical findings. As an ongoing project, we successfully construct higher order efficient positivity preserving time-stepping schemes for the Aït-Sahalia model [11] and try to cover a larger range of stochastic models [15].

### v2

@@ 1. Case 1:
    $\kappa+1>2\rho$. Noticing that $\gamma$ is an arbitrary large constant ensuring $\kappa\leq\gamma+\tfrac{1}{2}$ in the non-critical case due to Assumption 3.3, with the aid of Lemma 2.2 and Lemma 3.6, one can infer that
@@
    Equipped with (4.12) as well as (4.3), Lemmas 2.4 and 2.6 one can directly deduce by the Jensen inequality and the Hölder inequality that
    $\mathbb{E}\Big{[}\big{|}R_{n+1}^{(1)}\big{|}^{2}\Big{]}$ $\leq 4\mathbb{E}\bigg{[}\Big{|}\int_{t_{n}}^{t_{n+1}}c_{-1}(X_{s}^{-1}-X_{t_{n+1}}^{-1})\text{d}s\Big{|}^{2}\bigg{]}+4\mathbb{E}\bigg{[}\Big{|}\int_{t_{n}}^{t_{n+1}}c_{1}(X_{s}-X_{t_{n}})\text{d}s\Big{|}^{2}\bigg{]}$
    $\quad+4\mathbb{E}\bigg{[}\Big{|}\int_{t_{n}}^{t_{n+1}}(f(X_{s})-f(X_{t_{n}}))\text{d}s\Big{|}^{2}\bigg{]}+4\mathbb{E}\bigg{[}\Big{|}\int_{t_{n}}^{t_{n+1}}(f(Y_{n})-f_{h}(Y_{n}))\text{d}s\Big{|}^{2}\bigg{]}$
    $\leq 4h\int_{t_{n}}^{t_{n+1}}\mathbb{E}\Big{[}\big{|}c_{-1}(X_{s}^{-1}-X_{t_{n+1}}^{-1})\big{|}^{2}\Big{]}\text{d}s+4h\int_{t_{n}}^{t_{n+1}}\mathbb{E}\Big{[}\big{|}c_{1}(X_{s}-X_{t_{n}})\big{|}^{2}\Big{]}\text{d}s$
    $\quad+4h\int_{t_{n}}^{t_{n+1}}\mathbb{E}\Big{[}\big{|}(f(X_{s})-f(X_{t_{n}}))\big{|}^{2}\Big{]}\text{d}s+4h\int_{t_{n}}^{t_{n+1}}\mathbb{E}\Big{[}\big{|}(f(Y_{n})-f_{h}(Y_{n}))\big{|}^{2}\Big{]}\text{d}s$
    $\leq Ch^{3}.$
    In a similar way, but with the Itô isometry leads to
    $\mathbb{E}\big{[}\big{|}R_{n+1}^{(2)}\big{|}^{2}\big{]}\leq Ch^{2}.$ (4.13)
@@ 2. Case 2:
    The difference for the critical case comes from the extra restriction that $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$, as required by Assumption 3.3, which together with Lemmas 2.3 and 3.6 ensures the moment boundedness of $\mathbb{E}\big{[}|X_{t}|^{p}\big{]}$ and $\mathbb{E}\big{[}|Y_{n}|^{p}\big{]}$ for $p\leq 2\gamma+1\leq\tfrac{2c_{2}}{c_{3}^{2}}+1$. Therefore, on the condition $\tfrac{c_{2}}{c_{3}^{2}}>\kappa-\tfrac{1}{2}$, one can derive $2\kappa$-th moment bounds for both analytic and numerical solutions
@@
    Recalling (3.8), one knows that
@@
    by Corollary 3.7. This together with Lemmas 2.5 and 2.6 similarly helps us to deduce that, on the more strict assumption $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$,
@@
Let $\{X_{t}\}_{0\leq t\leq T}$ and $\{Y_{n}\}_{0\leq n\leq N}$ be defined as (2.2) and (3.1), respectively. Let $f_{h}$ and $g_{h}$ be defined as in Example 3.5. Suppose that Assumptions 3.2, 3.3, and 3.4 hold, with additionally one of the following conditions standing:
@@
Then, there exists a constant $C>0$ independent of the step-size $h$ such that
@@
We would like to mention that, for the general critical case $\kappa+1=2\rho$, one can observe a slightly more strict restriction on the model parameters than that for the implicit methods (STMs) in [22], where $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$ is required. This is caused by the need of $m_{1},m_{2}\leq 2\gamma+1$ in order to confirm Assumption 3.4. More precisely, as indicated by Example 3.5, for the general critical case, $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ as well as $12{\color[rgb]{0,0,0}(2\alpha\kappa+\rho)}=m_{2}<m_{1}=2(2\alpha+1)\kappa\leq 2\gamma+$ 1 are required, leading to $(2\alpha+1)\kappa\leq\tfrac{c_{2}}{c_{3}^{2}}+\tfrac{1}{2}$ as mentioned.
@@ 5 Numerical experiments
The semi-implicit tamed Euler-type method (TEM) in Example 3.5 with $\alpha=\tfrac{1}{2}$ applied to the model reads:
@@
The following three sets of parameters are taken as numerical tests.
@@
${\color[rgb]{0,0,0}\kappa=3,\rho=2},c_{-1}=1.5,c_{0}=2,c_{1}=1,c_{2}=4,{\color[rgb]{0,0,0}c_{3}=0.5}$.
Example 5.3.
$\kappa=2,\rho=1.5,c_{-1}=2,c_{0}=3,c_{1}=4,c_{2}=7,c_{3}=\sqrt{2}$.
Note that Example 5.1 satisfies the non-critical condition $\kappa+1>2\rho$, Example 5.2 satisfies the critical condition $\kappa+1=2\rho$, and Example 5.3, borrowed from [10, SET II], satisfies the critical condition with $\kappa=2$, $\rho=1.5$ and $\tfrac{c_{2}}{c_{3}^{2}}=2\kappa-\tfrac{1}{2}$. The strong convergence rates are tested with the endpoint $T=1$. For comparison, the backward Euler method (BEM) [20], which has been proved to be strong convergent with order 0.5 for the model [22], is also tested. Moreover, numerical approximations produced by the BEM with the step-size $h_{exact}=2^{-14}$ is used to simulate the “exact” solutions, while $h=2^{-i},i=4,5,...,9$ are used for numerical solutions over $10^{4}$ Brownian paths. The estimation error is calculated in terms of $e_{h}:=\bigg{(}\sup\limits_{0\leq n\leq N}\mathbb{E}\Big{[}|X_{t_{n}}-Y_{n}|^{2}\Big{]}\bigg{)}^{\tfrac{1}{2}}$.
Fig. 1 and Fig. 2 plot the average sample errors against various step-sizes on a log-log scale for Examples 5.1, 5.2, and 5.3, with BEM in a blue solid line, TEM in a red solid line, and the boundary preserving explicit scheme (BPES) proposed in [10] in a green solid line. The expected order $0.5$ of strong convergence is easily detected for both BEM and TEM. By a least square fitting, the resulting convergence rate $q$ and the least squares residual are shown in Table 1, where one can also confirm the predicted convergence rate. For Example 5.3 of the particular critical case $\kappa=2,\rho=1.5$, the BPES shows order one convergence and outperforms the other two order one-half schemes in terms of accuracy. Nevertheless, the BPES is simply designed for the particular critical case $\kappa=2,\rho=1.5$ (see [10] for more details), but the proposed explicit scheme, TEM, works for all cases.
In addition, the computational costs of BEM and the proposed TEM scheme over $10^{4}$ Brownian paths are shown in Table 2, and 50 out of the $10^{4}$ paths are shown in Fig. 3. It is easy to see that the TEM can preserve positivity and reduce computational costs significantly, compared with BEM, especially as the parameter $\kappa$ increases.
Figure 1: Strong convergence rate of TEM and BEM for Example 5.1 (left) and Example 5.2 (right).
Figure 2: Strong convergence rate of TEM and BEM for Example 5.3.
Figure 3: Fifty paths of numerical solutions for BEM and TEM
Table 1: A least square fit for the convergence rate $q$
@@
      Eg. 5.1       $q=0.5409$, $resid=0.0269$        $q=0.4458$, $resid=0.0074$
      Eg. 5.2        $q=0.6658$, $resid=0.1161$       $q=0.5347$, $resid=0.0524$
      Eg. 5.3       $q=0.5095$, $resid=0.0352$       $q=0.4906$, $resid=0.0174$
@@ 6 Conclusion
In this paper, we propose and analyze an explicit time-stepping scheme for the strong approximations of the generalized Aït-Sahalia model. Based on a semi-implicit strategy and proper modifications, the numerical scheme is designed to be positivity preserving and easily implementable with cheap computational costs compared with existing positivity preserving schemes. The mean-square strong convergence rate of the scheme is carefully analyzed and proved to be order $0.5$. Numerical experiments are also provided to support the theoretical findings. As an ongoing project, we successfully construct higher order efficient positivity preserving time-stepping schemes for the Aït-Sahalia model [12] and try to cover a larger range of stochastic models [17].